OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x - 2y = 6. Answer x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (aravindg):

divide whole equation by 6

OpenStudy (aravindg):

u will get \[\large \frac{x}{2}+\frac{y}{-3}=1\] this is intercept form where a=2 and b=-3 so coordinates are (2,-3)

OpenStudy (aravindg):

understood?

OpenStudy (auctoratrox):

you can put this in slope-intercept form. you get y = (3/2)x -3, so inputing zero for both x and y gives (2,0) and (0,-3)

OpenStudy (anonymous):

to get x intercept change y to 0, to get y intercept change x to 0 3x - 2y = 6 3(0)-2y =6 -2y=6 y=-3 (0,-3) y-intercept 3x - 2y = 6 3x-2(0) = 6 3x=6 x=2 (2,0) x intercept